Wilmington Company has two manufacturing departments—Assembly and Fabrication. It considers all of its manufacturing overhead costs to be fixed costs. The first set of data that is shown below is based on estimates from the beginning of the year. The second set of data relates to one particular job completed during the year—Job Bravo.

Estimated Data Assembly Fabrication Total
Manufacturing overhead costs $ 4,370,000 $ 4,830,000 $ 9,200,000
Direct labor-hours 115,000 69,000 184,000
Machine-hours 46,000 230,000 276,000
Job Bravo Assembly Fabrication Total
Direct labor-hours 24 16 40
Machine-hours 16 19 35

Required:

1. If Wilmington used a plantwide predetermined overhead rate based on direct labor-hours, how much manufacturing overhead would be applied to Job Bravo?

2. If Wilmington uses departmental predetermined overhead rates with direct labor-hours as the allocation base in Assembly and machine-hours as the allocation base in Fabrication, how much manufacturing overhead would be applied to Job Bravo?

(Round your intermediate calculations to 2 decimal places.)
Answer #1

1) Overhead rate = 9200000/184000 = 50 per DLH

Overhead applied to Job Bravo = 40*50 = $2000

2) Assembly overhead rate = 4370000/115000 = 38 per DLH

Fabrication overhead rate = 4830000/230000 = 21 Per MH

Overhead applied to Job Bravo = (24*38+19*21) = 1311